# Credibility-Aware Multi-Modal Fusion Using Probabilistic Circuits

### v1

@@ 1 Introduction
Real-world decision-making in high-profile tasks such as healthcare requires learning and reasoning reliably by utilizing the diverse modalities of available data sources. While such multi-modal data offers a rich representation and potentially multiple views of the underlying phenomena (for example, images vs blood tests in a clinical setting), it also makes learning and inference more challenging. Raw data from different sources is often noisy, incomplete, and inconsistent. This heterogeneity poses a significant obstacle to effective data fusion and analysis.
@@
Multi-modal fusion techniques [Baltrušaitis et al., 2018] have emerged as a promising approach to combine information from multiple sources to enhance performance on discriminative learning tasks. These techniques aim to extract and integrate complementary information from different modalities, leading to more robust and reliable results. However, a crucial aspect that often remains overlooked in multimodal fusion is explicit modeling of the credibility of the information sources. In many applications, such as sensor fusion [Khaleghi et al., 2013], medical diagnosis [Kline et al., 2022], and financial analysis [Sawhney et al., 2020], the quality and reliability of the information sources vary significantly. Distinguishing reliable sources from non-reliable sources is essential for making accurate and informed decisions. Multimodal fusion methods often assume that all sources are equally credible, which can lead to suboptimal performance or even erroneous conclusions.
Credibility-aware methods in the context of late multimodal fusion have previously used weighted average [Rogova and Nimier, 2004], discounting factors [Elouedi et al., 2004b] and Bayesian networks [Wright and Laskey, 2006]. This results in models of credibility that are either too simple (as in the case of weighted averages and discounting factors) to model complex dependencies or too complex to perform tractable inference (as in the case of general Bayesian networks or more recent deep models).
@@
We focus on multi-modal discriminative learning and propose a late fusion method that uses Probabilistic Circuits (PCs) [Choi et al., 2020], to effectively combine the predictive distributions over individual modalities. PCs are a class of generative models that are expressive enough to model complex distributions while tractable for exact inference. Using the tractability of PCs, we define a probabilistic measure for assessing the credibility. Some salient features of our approach are that the use of PCs (1) allows for modeling uncertainty over unimodal predictive distributions effectively; (2) makes the model robust to noise and outliers; (3) enables effective handling of missing data; (4) is grounded in a robust theoretical framework; and (4) finally, makes it possible to obtain faithful estimates of credibility.
@@
Our paper makes the following key contributions: (1) As far as we are aware, we introduce the first theoretically grounded multimodal fusion with strong probabilistic semantics based on PCs; (2) We present two versions of our late fusion algorithm with different characteristics; (3) We derive a theoretically grounded measure of credibility and illustrate its connection to the conditional entropy over unimodal predictive distributions, allowing for reliable late fusion; (4) Finally, we experimentally validate the efficacy of PCs in modeling complex interactions between modalities and faithfully estimating their credibility.
@@
The rest of the paper is organized as follows: we begin with a concise overview of essential background and relevant works. Following this, we formulate the problem at hand and our PC-based fusion method, along with the architectural details and methodology for assessing credibility. We then experimentally evaluate the effectiveness of our method and finally conclude by summarizing our findings, contributions, and future work.
@@ 2.2 Credibility
We follow prior works that approach the problem of accounting for source reliability in multimodal fusion from the perspective of the credibility of the information provided by the source. These works perform multimodal fusion using source-reliability coefficients learned using domain and contextual information  [Nimier, 1998, Fabre et al., 2001]. In the absence of such information, an alternate approach involves learning these coefficients from training data. This is achieved by minimizing the distance between a vector of beliefs resulting from fusion and a target vector from the training set  [Rogova and Kasturi, 2001, Elouedi et al., 2004a]. Another method for establishing reliability, by using training data, is based on separability, wherein the average statistical separability of information classes in each source is considered [Benediktsson et al., 1990]. This category of methods i.e. learning coefficients from training data, proves useful in establishing the relative credibility of the predictions of classifiers.
@@ 2.3 Probabilistic circuits (PCs)
$\small P_{n}(\mathbf{X}=\mathbf{x})=\begin{cases}\sum_{c\in\textbf{ch}(n)}w_{c}P_{c}(\mathbf{X}=\mathbf{x})&n\in\text{Sum}\\ \prod_{c\in\textbf{ch}(n)}P_{c}(\mathbf{X}_{\textbf{sc}(c)}=\mathbf{x}_{\textbf{sc}(c)})&n\in\text{ Product}\\ \psi_{n}(\mathbf{X}=\mathbf{x})&n\in\text{Leaf}\\ \end{cases}$
@@
The key advantage of PCs is that they admit tractable and often linear time inference for a variety of probabilistic queries under mild assumptions about the structure of $G$. In this work, we consider a subclass of PCs that are smooth and decomposable (typically called sum-product networks  [Poon and Domingos, 2011]). A PC satisfies smoothness if the scope of each sum node is identical to the scope of each of its children. It satisfies decomposability if, for each product node, all the children have disjoint scopes. Smoothness and decomposability allow us to tractably infer marginal and conditional distributions from the learned joint.
@@ 3 Multimodal fusion via PCs
We begin by formalizing the noisy late multi-modal fusion setting for discriminative learning that we focus on in this work. Given a dataset in which features predictive of a target concept are obtained from multiple different modalities, the late fusion setting involves training an expert over each modality to estimate the unimodal predictive distribution over the target and then combining them using a fusion function (probabilistic combination function in our case) to obtain the final output. More formally,
Given: A dataset $\mathcal{D}=\{(\mathbf{x}_{1}^{i},\mathbf{x}_{2}^{i}\ldots\mathbf{x}_{M}^{i},y^{i})\}_{i=1}^{N}$ with $N$ data points, each with information from $M$ different modalities, i.e. each $\mathbf{x}^{i}_{j}\in\mathbb{R}^{d_{j}}$ where $d_{j}$ denotes the feature dimension corresponding to modality $j$ for the $i^{th}$ example, and $y^{i}$ is its target.
@@
We now outline more formally how the defined notion of credibility is related to the uncertainty over the unimodal predictive distribution. A well-established method for quantifying the uncertainty and information content within a random variable is through the concept of entropy. We establish the following theorem that correlates the credibility of a modality with the entropy of its predictive distribution.
@@
The expected credibility $\mathcal{C}^{j}$ of a modality $j$ in predicting the target $Y$, is lower bounded by the negative of the conditional entropy $(\mathbb{H})$ of the unimodal predictive distribution of modality $j$ over $Y$, given the predictive distributions of all other modalities, i.e.
@@
$\mathcal{C}^{j}={}$ $\delta(P(Y|\mathbf{F})||P(Y|\mathbf{F}^{-j}))$
[…]
Now, we know that $\log\frac{P(\mathbf{F}^{-j})}{P(\mathbf{F})}\geq 0$ as $P$ is a probability mass function and $P(x_{1},x_{2})\geq 0,\quad\forall x_{1},x_{2}\implies\sum_{x_{2}}P(x_{1},x_{2})\geq P(x_{1},x_{2})$. Thus, we have
[…]
$\mathcal{C}^{j}\geq{}$ $-\mathbb{H}(\mathcal{F}_{\phi_{j}}|\mathbf{F}^{-j})$
Intuitively, a modality that is less corrupted by noise and more informative of the target than others can be expected to have a lower predictive entropy. Thus, by the above theorem, we can conclude that such a modality would always have a higher assigned credibility than others. Conversely, when a modality gets corrupted by noise, its credibility score decreases. Thus the defined measure of credibility is theoretically grounded. Its utility becomes evident in critical domains such as healthcare, where the stakes of decision-making are high. In such contexts, credibility assessments can guide the reliance on specific expert systems or enable the discounting of modalities that are deemed unreliable.
@@ 3.1 PCs as combination functions
We now present the details of late fusion models $\mathcal{M}$ capable of incorporating the above-defined notion of credibility. It is clear that estimating credibility requires access to a generative model that estimates the joint distribution over $Y$ and the unimodal predictors $\{\mathcal{F}_{j}\}_{j=1}^{M}$. Additionally, the generative model should support efficient and exact evaluation of both joint and conditional probability densities. Probabilistic Circuits (PCs) are one such class of generative models that can model complex distributions while supporting tractable and linear time inference of conditional and marginal distributions.
@@
Given a multi-modal example, $(\mathbf{x}_{1},\dots,\mathbf{x}_{M}),$ we can perform predictive inference over target $Y$ as follows:
  1. 1.
    Compute $\mathbf{p}_{j}=\mathcal{M}_{\phi_{j}}(\mathbf{x}_{j})$ for each modality $j$ by evaluating the unimodal predictors.
  2. 2.
    Infer the multimodal predictive distribution over $Y$ given the unimodal distributions $\mathbf{p}_{1},\dots,\mathbf{p}_{M}$ by performing conditional inference: $P_{\theta}(Y\mid\mathbf{p}_{1},\dots,\mathbf{p}_{M})=\frac{P_{\theta}(Y,\mathbf{p}_{1},\dots,\mathbf{p}_{M})}{P_{\theta}(\mathbf{p}_{1},\dots,\mathbf{p}_{M})}$
The credibility of a modality $j$ can then be estimated using the PC $\theta$ as $\mathcal{C}^{\theta}_{j}=\delta(P_{\theta}(Y|\mathbf{p}_{1},\ldots\mathbf{p}_{M})||P_{\theta}(Y|\mathbf{p}_{1},\ldots\mathbf{p}_{j-1},\mathbf{p}_{j+1}\ldots\mathbf{p}_{M}))$
An alternative to the Direct-PC combination function, that explicitly utilizes the credibility scores would be to define the final predictive distribution as a convex sum of credibility-weighted unimodal predictive distributions. i.e:
@@
initialize: $\tilde{\theta}=\theta,\{\tilde{\phi_{j}}=\phi_{j}\}_{j=1}^{M},t=1$ while $t\leq t_{\text{max}}$ do
       $\{(\mathbf{x}_{j}^{i},y^{i})_{j=1}^{M}\}_{i=1}^{B}\sim\mathcal{D}$triangleright Sample a mini-batch For each modality $j$ and data point $i$
       $\triangleright$ Compute unimodal predictive distributions ${\mathbf{p}^{i}_{j}}$ ${\mathbf{p}_{j}^{i}}\leftarrow\mathcal{M}_{\tilde{\phi_{j}}}(\mathbf{x}_{j}^{i})$
       $\triangleright$ Obtain credibility scores $\mathcal{C}_{j}^{i}\leftarrow\delta(P_{\tilde{\theta}}(Y|\{\mathbf{p}^{i}_{k}\}_{k=1}^{M})||P_{\tilde{\theta}}(Y|\{\mathbf{p}^{i}_{k}\}_{k=1}^{M}\setminus\mathbf{p}_{j}^{i}))$ $\tilde{\mathcal{C}}_{j}^{i}\leftarrow\mathcal{C}_{j}^{i}/(\sum_{j=1}^{M}\mathcal{C}_{j}^{i})$
       $\triangleright$ Compute the final predictive distribution $\mathbf{p}^{i}\leftarrow\sum_{j=1}^{M}\tilde{\mathcal{C}}_{j}^{i}\mathbf{p}_{j}^{i}$ if CWM else $P_{\tilde{\theta}}(Y|\{\mathbf{p}^{i}_{k}\}_{k=1}^{M})$
       $\triangleright$ Compute the empirical loss $L_{j}\leftarrow\frac{1}{B}\sum_{i=1}^{B}l(\mathbf{p}_{j}^{i},y^{i})$ $L\leftarrow\frac{1}{B}\sum_{i=1}^{B}l(\mathbf{p}^{i},y^{i})+\sum_{j=1}^{M}L_{j}$
       $\triangleright$ Update the unimodal predictors and PC $\{\tilde{\phi_{j}}\}_{j=1}^{M}\leftarrow\{\tilde{\phi_{j}}\}_{j=1}^{M}-\eta_{1}\nabla_{\{\tilde{\phi_{j}}\}_{j=1}^{M}}L$ $\tilde{\theta}\leftarrow\tilde{\theta}-\eta_{2}\nabla_{\tilde{\theta}}L+\eta_{2}\nabla_{\tilde{\theta}}\sum_{i=1}^{B}P_{\tilde{\theta}}(y^{i},\{\mathbf{p}_{j}^{i}\}_{j=1}^{M})$ $t=t+1$

### v2

@@ 1 Introduction
Real-world decision-making in high-profile tasks such as healthcare requires learning and reasoning reliably by utilizing the diverse modalities of available data sources. While such multi-modal data offer rich representations and potentially multiple views of the underlying phenomena (for example, images vs blood tests in a clinical setting), they also make learning and inference quite challenging. Raw data from different sources is often noisy, incomplete, and inconsistent. This heterogeneity poses a significant obstacle to effective data fusion and analysis.
@@
Multi-modal fusion techniques [Baltrušaitis et al., 2018] have emerged as a promising approach to combine information from multiple sources to enhance performance on discriminative learning tasks. These techniques aim to extract and integrate complementary information from different modalities, leading to more robust and reliable results. However, a crucial aspect that often remains overlooked in multimodal fusion is the explicit modeling of the credibility of the information sources. In many applications, such as sensor fusion Khaleghi et al. [2013], medical diagnosis Kline et al. [2022], and financial analysis Sawhney et al. [2020], the quality and reliability of the information sources vary significantly. Distinguishing reliable sources from non-reliable sources is essential for making accurate and informed decisions. Multimodal fusion methods often assume that all sources are equally credible, which can lead to suboptimal performance or even erroneous conclusions.
@@
Credibility-aware methods in the context of late multimodal fusion have previously used weighted average Rogova and Nimier [2004], discounting factors Elouedi et al. [2004b] and Bayesian networks Wright and Laskey [2006]. This results in models of credibility that are either too simple (as in the case of weighted averages and discounting factors) to model complex dependencies or too complex to perform tractable inference/reasoning (as in the case of general Bayesian networks or more recent deep models). We focus on multi-modal discriminative learning and propose a late fusion method that uses Probabilistic Circuits (PCs) Choi et al. [2020], to effectively combine the predictive distributions over individual modalities. PCs are a class of generative models that are expressive enough to model complex distributions while tractable for exact inference. Using the tractability of PCs, we define a probabilistic measure for assessing the credibility. Some salient features of our approach are that the use of PCs (1) allows for modeling uncertainty over unimodal predictive distributions effectively; (2) makes the model robust to noise and outliers; (3) enables effective handling of missing data; (4) is grounded in a robust theoretical framework; and (4) finally, makes it possible to obtain faithful estimates of credibility.
@@
Our paper makes the following key contributions: (1) To our knowledge, we introduce the first theoretically grounded multimodal fusion with strong probabilistic semantics based on PCs; Specifically, we identify the class of PCs that are amenable to this task of credibility-aware multi-modal fusion and define their characteristics; (2) We present two versions of our late fusion algorithm with different characteristics; (3) We derive a theoretically grounded measure of credibility and illustrate its connection to the conditional entropy over unimodal predictive distributions, allowing for reliable late fusion; (4) Finally, we experimentally validate the efficacy of PCs in modeling complex interactions between modalities and faithfully estimating their credibility.
@@
The rest of the paper is organized as follows: we begin with a concise overview of essential background and relevant work. Following this, we formulate the problem at hand and our PC-based fusion method, along with the architectural details and methodology for assessing credibility. We then experimentally evaluate the effectiveness of our method and finally conclude by summarizing our findings, contributions, and future work.
@@ 2.2 Credibility
We follow prior works that approach the problem of accounting for source reliability in multimodal fusion from the perspective of the credibility of the information provided by the source. These works perform multimodal fusion using source-reliability coefficients learned using domain and contextual information  [Nimier, 1998, Fabre et al., 2001]. In the absence of such information, an alternate approach involves learning these coefficients from data. This is achieved by minimizing the distance between a vector of beliefs resulting from fusion and a target vector from the training set  [Rogova and Kasturi, 2001, Elouedi et al., 2004a]. Another data-driven method for establishing reliability is based on separability, wherein the average statistical separability of information classes in each source is considered [Benediktsson et al., 1990]. This category of methods i.e. learning coefficients from training data, proves useful in establishing the relative credibility of the predictions of classifiers.
@@ 2.3 Probabilistic Circuits (PCs)
$\small P_{n}(\mathbf{X}=\mathbf{x})=\begin{cases}\sum_{c\in\textbf{ch}(n)}w_{c}P_{c}(\mathbf{X}=\mathbf{x})&n\in\text{Sum}\\ \prod_{c\in\textbf{ch}(n)}P_{c}(\mathbf{X}^{\textbf{sc}(c)}=\mathbf{x}^{\textbf{sc}(c)})&n\in\text{ Product}\\ \psi_{n}(\mathbf{X}=\mathbf{x})&n\in\text{Leaf}\\ \end{cases}$
@@
The key advantage of PCs is that they admit tractable and often linear time inference for a variety of probabilistic queries under mild assumptions about the structure of $G$. In this work, we consider a subclass of PCs that are smooth and decomposable (typically called sum-product networks  Poon and Domingos [2011]). A PC satisfies smoothness if the scope of each sum node is identical to the scope of each of its children. It satisfies decomposability if, for each product node, all the children have disjoint scopes. Smoothness and decomposability allow us to tractably infer marginal and conditional distributions from the learned joint.
@@ 3 Multimodal fusion via PCs
We begin by formalizing the noisy late multi-modal fusion setting for discriminative learning that we focus on. Given a dataset in which features predictive of a target concept are obtained from multiple different modalities, the late fusion setting involves training an expert over each modality to estimate the unimodal predictive distribution over the target and then combining them using a fusion function (probabilistic combination function in our case) to obtain the final output. More formally,
Given: A dataset $\mathcal{D}=\{(\mathbf{x}_{1}^{i},\mathbf{x}_{2}^{i}\ldots\mathbf{x}_{M}^{i},y^{i})\}_{i=1}^{N}$ with $N$ data points, each with information from $M$ different modalities, i.e. each $\mathbf{x}^{i}_{j}\in\mathbb{R}^{d_{j}}$ where $d_{j}$ denotes the feature dimension corresponding to modality $j$ for the $i^{th}$ example, and $y^{i}$ denotes its target class.
@@
We now outline more formally how the defined notion of credibility is related to the uncertainty over the unimodal predictive distribution. A well-established method for quantifying the uncertainty and information content within a random variable is through the concept of entropy. We present a theorem that correlates the credibility of a modality with the entropy of its predictive distribution, under mild assumptions, as defined below.
Definition 2.
A model $\mathcal{M}$ representing a probability distribution ($P_{\mathcal{M}}$) over $n$ random variables $\mathbf{X}$ is said to be Marginal Dominant if its marginals are lower bounded by the joint everywhere. i.e.,
$P_{\mathcal{M}}(\mathbf{X}^{-j}=\mathbf{x}^{-j})$ $\geq P_{\mathcal{M}}(\mathbf{X}^{-j}=\mathbf{x}^{-j},\mathbf{X}^{j}=\mathbf{x}^{j})$
$\forall\ (\mathbf{x}^{j},\mathbf{x}^{-j})\in Dom.(\mathbf{X}^{j},\mathbf{X}^{-j})$
where $\ j\subseteq\{1,\dots,n\}$ and we use the notation $\mathbf{X}^{-j}$ to denote $\{X_{i}\}_{i=1}^{n}\setminus\{X_{k}\}_{k\in j}$ and $Dom.(\mathbf{X})$ to denote the domain set of the variables in $\mathbf{X}$.
@@
The expected credibility $\mathcal{C}^{j}$ of a modality $j$ in predicting the target $Y,$ under a Marginal Dominant distribution is lower bounded by the negative of the conditional entropy $(\mathbb{H})$ of the unimodal predictive distribution of modality $j$ over $Y,$ given the predictive distributions of all other modalities, i.e.
@@
Deferred to the appendix. ∎
Intuitively, a modality less corrupted by noise and more informative of the target than others can be expected to have a lower predictive entropy. Thus, by the above theorem, we can conclude that such a modality would always have a higher assigned credibility than others. Conversely, when a modality gets corrupted by noise, its credibility score decreases. Thus the defined measure of credibility is theoretically grounded. Its utility becomes evident in critical domains such as healthcare, where the stakes of decision-making are high. In such contexts, credibility assessments can guide the reliance on specific expert systems or enable the discounting of modalities that are deemed unreliable.
@@ 3.1 PCs as Combination Functions
We now present the details of late fusion models $\mathcal{M}$ capable of incorporating the above-defined notion of credibility. It is clear that estimating credibility requires access to a generative model that estimates the joint distribution over $Y$ and the unimodal predictors $\{\mathcal{F}_{j}\}_{j=1}^{M}$. Additionally, the generative model should support efficient and exact evaluation of both joint and conditional probability densities. Probabilistic Circuits (PCs) are one such class of generative models that can model complex distributions while supporting tractable and linear time inference of conditional and marginal distributions. Further, as we show below, the distribution modeled by a PC is Marginal Dominant under certain structural properties, making it well-suited for credibility-aware fusion.
Theorem 3.2.
A Probabilistic Circuit is Marginal Dominant  if it is smooth, decomposable, and has leaf distributions with unimodal densities upper-bounded by unity.
Proof.
Deferred to the appendix. ∎
@@
Predictive Inference. Given a multi-modal example, $(\mathbf{x}_{1},\dots,\mathbf{x}_{M}),$ we can perform predictive inference over target $Y$ as follows:
(1) compute $\mathbf{p}_{j}=\mathcal{M}_{\phi_{j}}(\mathbf{x}_{j})$ for each modality $j$ by evaluating the unimodal predictors $\mathcal{M}_{i},\dots,\mathcal{M}_{m}$. (2) Infer the multimodal predictive distribution over $Y$ given the unimodal distributions $\mathbf{p}_{1},\dots,\mathbf{p}_{M}$ by performing conditional inference:
$P_{\theta}(Y\mid\mathbf{p}_{1},\dots,\mathbf{p}_{M})$ $=\frac{P_{\theta}(Y,\mathbf{p}_{1},\dots,\mathbf{p}_{M})}{P_{\theta}(\mathbf{p}_{1},\dots,\mathbf{p}_{M})}$
$=\frac{P_{\theta}(Y,\mathbf{p}_{1},\dots,\mathbf{p}_{M})}{\sum_{y}P_{\theta}(Y=y,\mathbf{p}_{1},\dots,\mathbf{p}_{M})}$
Credibility Assessment. The credibility of a modality $j$ can then be estimated using the PC $\theta$ as
$\mathcal{C}^{\theta}_{j}=\delta(P_{\theta}(Y|\mathbf{p}_{1},\ldots\mathbf{p}_{M})||P_{\theta}(Y|\mathbf{p}_{1},\ldots\mathbf{p}_{j-1},\mathbf{p}_{j+1}\ldots\mathbf{p}_{M}))$
As smooth and decomposable PCs support linear time evaluation of joint, marginal, and conditional distributions, both predictive and credibility inference can thus be achieved in linear time.
An alternative to the Direct-PC combination function, which explicitly utilizes the credibility scores would be to define the final predictive distribution as a convex sum of credibility-weighted unimodal predictive distributions. i.e:
@@
initialize: $\tilde{\theta}=\theta,\{\tilde{\phi_{j}}=\phi_{j}\}_{j=1}^{M},t=1$
while $t\leq t_{\text{max}}$ do
      $\{(\mathbf{x}_{j}^{i},y^{i})_{j=1}^{M}\}_{i=1}^{B}\sim\mathcal{D}$triangleright Sample a mini-batch
      For each modality $j$ and data point $i$
       $\triangleright$ Compute unimodal predictive distributions ${\mathbf{p}^{i}_{j}}$
      ${\mathbf{p}_{j}^{i}}\leftarrow\mathcal{M}_{\tilde{\phi_{j}}}(\mathbf{x}_{j}^{i})$
       $\triangleright$ Obtain credibility scores
      $\mathcal{C}_{j}^{i}\leftarrow\delta(P_{\tilde{\theta}}(Y|\{\mathbf{p}^{i}_{k}\}_{k=1}^{M})||P_{\tilde{\theta}}(Y|\{\mathbf{p}^{i}_{k}\}_{k=1}^{M}\setminus\mathbf{p}_{j}^{i}))$
      $\tilde{\mathcal{C}}_{j}^{i}\leftarrow\mathcal{C}_{j}^{i}/(\sum_{j=1}^{M}\mathcal{C}_{j}^{i})$
       $\triangleright$ Compute the final predictive distribution
      $\mathbf{p}^{i}\leftarrow\sum_{j=1}^{M}\tilde{\mathcal{C}}_{j}^{i}\mathbf{p}_{j}^{i}$ if CWM else $P_{\tilde{\theta}}(Y|\{\mathbf{p}^{i}_{k}\}_{k=1}^{M})$
       $\triangleright$ Compute the empirical loss
      $L_{j}\leftarrow\frac{1}{B}\sum_{i=1}^{B}l(\mathbf{p}_{j}^{i},y^{i})$
      $L\leftarrow\frac{1}{B}\sum_{i=1}^{B}l(\mathbf{p}^{i},y^{i})+\sum_{j=1}^{M}L_{j}$
       $\triangleright$ Update the unimodal predictors and PC
      $\{\tilde{\phi_{j}}\}_{j=1}^{M}\leftarrow\{\tilde{\phi_{j}}\}_{j=1}^{M}-\eta_{1}\nabla_{\{\tilde{\phi_{j}}\}_{j=1}^{M}}L$
      $\tilde{\theta}\leftarrow\tilde{\theta}-\eta_{2}\nabla_{\tilde{\theta}}L+\eta_{2}\nabla_{\tilde{\theta}}\sum_{i=1}^{B}P_{\tilde{\theta}}(y^{i},\{\mathbf{p}_{j}^{i}\}_{j=1}^{M})$
      $t=t+1$